Circular No. 6442

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1996an IN NGC 1084 S. Nakano, Sumoto, Japan, reports the discovery by Masakatsu Aoki,

Tsukioka-cho, Toyama, of a supernova (mag 14) in NGC 1084, found on CCD frames taken on July 27.772 UT with a 0.43-m f/5 reflector. Y. Kushida, Yatsugatake South Base Observatory, provides the following precise position from Aoki’s exposure for SN 1996an: α = 2^{h}46^{m}00^{s}.83, δ = −7^{o}34′20′′.3 (equinox 2000.0), which is 15′′ east and 22′′ north of the center of NGC 1084. R. Kushida (Yatsugatake) reports that a frame taken on 1995 Dec. 21 shows no star at the position of SN 1996an. SN 1963P also appeared in NGC 1084.

S. Benetti, M. Turatto, and G. van de Steene, European Southern Observatory (ESO), report: “Inspection of a fully-reduced CCD spectrogram (range 400–980 nm, resolution 0.7 nm) obtained on July 30.41 UT with the ESO 1.5-m telescope (+ Boller & Chivens spectrograph) at La Silla confirms this object as a type-II supernova, ∼ 2 months after explosion. The spectrum consists of a multitude of lines with the canonical P-Cyg profiles, superimposed on a relatively red continuum. Most of the lines arises from H I, Ca II, Na I, Ba II, Fe II, O I, and Sc II. The expansion velocities deduced from the minima of the most prominent lines are: 4250 km/s (Hα); 4000 km/s (Ca II infrared); 3650 km/s (Na I); 2600 km/s (Ba II); and 3000 km/s (Fe II). The recession velocity at the supernova location is 1250 km/s, as deduced from the narrow Hα and [N II] emissions. A narrow interstellar absorption of Na I D (measured at 591.8 nm) is clearly visible.”

COMET 22P/KOPFF Visual m_{1} and coma-diameter estimates (cf. IAUC 6414): June 8.97

UT, 7.8, 7′ (K. Hornoch, Lelekovice, Czech Republic, 10×80 binoculars); 18.35, 8.5, 4′ (R. Keen, Mt. Thorodin, CO, 0.15-m reflector); July 9.91, 7.8, 10′ (M. V. Zanotta, Laino, Italy, 20×80 binoculars); 12.17, 7.6, 7′.2 (J. E. Bortle, Stormville, NY, 20×80 binoculars); 24.53, 8.0, 6′ (A. Pearce, Elwood, Vic., Australia, 20×80 binoculars); 27.28, 8.3, 6′.5 (C. E. Spratt, Victoria, BC, 0.20-m reflector).

COMET C/1996 J1 (EVANS-DRINKWATER) Total visual magnitude estimates (cf. IAUC 6418): June 19.17 UT, 13.3

(A. Hale, Cloudcroft, NM, 0.41-m reflector); July 12.80, 13.1 (H. Luethen, Tivoli, Namibia, 0.25-m reflector); 17.80, 12.9 (Luethen).

1996 July 30  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green